one number is 5 more than another number. three times the first plus twice the second is 30. find the numbers
Answer by checkley77(12844)

Let x & x+5 be the 2 numbers.
3x+2(x+5)=30
3x+2x+10=30
5x=30-10
5x=20
x=20/5
x=4 ans. for the smaller number.
4+5=9 ans. for the larger number.
Proof:
3*4+2*9=30
12+18=30
30=30